# Can Neural Decompilation Assist Vulnerability Prediction on Binary Code?

### v1

@@ 4.2. Evaluation - Neural Decompilation
We evaluated the translation quality using output similarity metrics including BLEU, METEOR, ROUGE, and Edit Distance (ED), illustrated in Table 3.5. Each metric expresses different characteristics of translation quality: BLEU allows capturing the precision of n-grams, giving importance to sequence length; ROUGE is useful in capturing code structure and overlapping semantic units, such as word sequences and key phrases; METEOR takes into account precision, completeness, and word order. However, metrics such as BLEU, originally developed for natural text translation, may not be particularly appropriate for evaluating translation between code sequences (tran2019does, ; liguori2023evaluates, ). ED is suitable for capturing small variations in the code and is representative of the number of changes to perform on the decompiled output. Normalized ED lies between 0 and 1 and for each $i_{th}$ sequence pair $(hyp_{i}-ref_{i})$ is calculated as: Normalized ED_i = 1 - ED(hypi,refi)max(len(refi),len(hypi)) Hence, a high Normalized ED value corresponds to an optimal translation. In the following, we refer to this metric as ED to maintain coherence with other works in this field.
Table 3.5 compares different decompilation solutions based on their approach, number of functions used for training and testing, maximum target length, and output similarity metrics. The results of neural decompilation are outstanding. Specifically, fairseq achieves a 59% ED score using about $75,000$ C/C++ functions for both training and testing with a maximum target length of 338 tokens. These results are close to those of other studies on this topic: Hosseini et al. (hosseini2022beyond, ) reached an 54% ED using about $2,000,000$ C/C++ functions with a maximum length of $271$, while Katz et al. (katz2018using, ) achieves an ED of 30% using about $700,000$ C-code snippets with a maximum target length of $88$. Our results are slightly higher than the former and significantly higher than the latter. It is worth noting that both studies leverage a number of samples at least an order of magnitude higher than ours. In addition, the function lengths analyzed with our approach are greater than those used by either work. For the sake of clarity, we note that these works adopt a complementary version of ED, less widespread for NMT tasks (liguori2023evaluates, ). We provide further details about these works in Section 6. Focusing on the other metrics obtained from fairseq, ROUGE-L shows that, in 77% of cases, the generated code contains word sequences similar to those of the reference. We observe a 4-gram overlap of 58% (BLEU-4) and a token order adherence of 64% (METEOR).
For the sake of completeness, we also evaluated the performance of a traditional decompiler, Ghidra (Ghidra, ). Specifically, we randomly sampled 1,000 programs from our decompiled test set. Ghidra relies on control flow graphs (CFGs), data flow graphs (DFGs), and other intermediate representations (IRs) to perform decompilation. Ghidra results, with a 23% ED, indicate significantly lower performance compared to studies utilizing neural network-based models, which suggests that traditional graph-based methods may struggle to capture the complexities of the source code as effectively as neural approaches. However, unlike a traditional decompiler, an NMT model always generates an answer despite having a low probability of approaching the correct one. In addition, one of the disadvantages of using a neural network is that there is less control over the translation, i.e., in case of errors it is very complex to determine which factor led the model to a specific decision over others. An interesting finding is that pre-trained models, i.e., CodeT5+ and CodeBERT, perform considerably worse than the model trained from scratch, i.e., fairseq. Above all, CodeBERT presents a discrepancy of around 30% on BLEU-4 and ROUGE compared to fairseq. In most of the metrics CodeT5+ scores about 10% lower than fairseq, reaching in some cases intermediate values between fairseq and CodeBERT, as in the case of the ED (49%). As noted by Tufano et al. (tufano2023automating, ), we can associate the negative impact of pre-training with the size of the training dataset. Pre-training helps boost performance only if the amount of fine-tuning data available is small. We deem our dataset with about 75,000 samples large enough to overwrite the parameters tuned in the pre-training phase. According to these results, we adopt fairseq as the neural decompiler for further analysis.
@@
*: the results on the multi-class classification only concern Flawfinder-compatible CWEs. °: the results on the multi-class classification concern different CWEs selected by Schaad et al. (schaad2023deep, ).
Binary Classification. The first analysis concerns the classification of the target code as vulnerable/non-vulnerable. We submitted the C/C++ code translated by the neural decompilation model to the classification process, adopting the same data pre-processing for each model. We selected accuracy and F1-Score as the metrics to evaluate the performance of the models.
Table LABEL:tab:Classifications shows the models’ performance for the binary classification on the decompiled C/C++ code. CodeGPT is the best-performing model with an accuracy and F1-Score of 95%. The performance is high for all models, with both metrics around 90%. The highest gap lies between SRNN and CodeGPT, exhibiting a discrepancy of about 7% in terms of F1-Score. Table LABEL:tab:Classifications also reports the performance of a state-of-the-art work on vulnerability prediction in binary files (schaad2023deep, ), which achieves an 88% F1-Score through the LLVM IR and the SRNN model, further described in Table 7 and Section 6.
Furthermore, we tested a static analysis tool, Flawfinder (flawfinder, ), on the same test set. This tool scans C/C++ source code and flags potential security holes. Flawfinder exhibits metrics that are close to 50% on average, which is equivalent to a random model. Therefore, the results support the implemented design choice of using DL models for vulnerability prediction instead of a static analysis tool. Finally, we leveraged Shap (NIPS2017_7062, ), a tool for explaining the output of our vulnerability prediction models, to better understand their behavior and which features are the most influential for the learning process, i.e., the tokens most exploited for classification. This is a helpful analysis to avoid falling into the so-called Clever Hans effect, which occurs when the model produces correct predictions based on ”wrong” features (kauffmann2020clever, ). The results revealed tokens in C/C++ code such as: if, new, strcat, funct, static, malloc, source, and printline. These results allow us to state that the model focuses on tokens that are relevant to implementing countermeasures to vulnerabilities, e.g., if constructs to check specific conditions (e.g., underflow or overflow), without taking decisions based on ”easier” alternatives, for instance, the length of the program under analysis (excluded by padding mechanism).
@@
Work Dataset Classes of Standardized Intermediate Vulnerability Binary Multi-class
Vulnerabilities Taxonomy Representation Predictor Classification Classification
Accuracy Accuracy
@@
Multi-class Classification. This task aims to identify the CWE associated with a vulnerable function. Specifically, the number of classes (CWEs) for this classification task is 20, plus an additional class associated with non-vulnerable (good) functions. The complexity of this task is related to the similarity between test cases of different classes. Therefore, a deep inspection of the test cases is necessary to determine the actual CWE to which they belong. We decided not to select classes based only on the highest number of occurrences in the dataset. Instead, we manually selected classes based on the dissimilarity between different CWEs, to reduce the confusion degree for the model towards similar CWEs. Consequently, to the best of our knowledge, we evaluated the difference between CWEs from a logical point of view. For instance, we selected one CWE between 190 Integer Overflow and 191 Integer Underflow because of their semantic similarity to prevent the situation in which the model hardly distinguished one from the other. Moreover, we discarded CWEs 546 Suspicious Comment and 615 Inclusion of Sensitive Information in Source Code Comments since all comments were removed during the pre-processing stage, described in Section 3.3. It is worth noting that each class needs a significant number of test cases so that the model can learn from as many examples as possible. The requirement on the number of tests is that each class has at least $300$ samples (at least $150$ bad and $150$ good examples). We prioritized the CWEs that belonged to the Top 25 CWEs published yearly by MITRE among the most represented ones, i.e., those with a reasonable amount of samples.
@@
Table LABEL:tab:Classifications shows the accuracy and F1-Score of the models for the multi-class classification task. The models’ performance decreased from the binary classification case, with CodeBERT performing the best with an accuracy of 83% and F1-Score of 82%. The lowest performance is again achieved by SRNN with 70% F1-Score, showing a discrepancy from CodeBERT of about 12%. The performance gap between our approach, utilizing CodeBERT, and the method proposed by Schaad et al. (schaad2023deep, ) is approximately 6% in terms of accuracy and 3% in the F1-score. However, it is important to note that the 20 classes analyzed in our study differ both in type and quantity from their 23 classes (schaad2023deep, ), as shown in Table 7.
@@
Qualitative analysis. We performed a qualitative analysis to provide deeper insight into the vulnerability prediction evaluation. Specifically, we refer to the CodeBERT model, i.e., the best-performing model in the multi-class prediction. Table 4.4 shows two vulnerable examples of multi-class classification. The first is an example of misprediction: the program associated with CWE 369 Divide By Zero was incorrectly associated with CWE 197 Numeric Truncation. The original program presented the instruction int result = (int)(100.0/data), i.e., the core of the vulnerability, completely omitted in the decompiled version. Despite a high ED metric, this error propagated into the prediction task, causing the model to fail.
The second example displays a correct prediction for CWE 190 Integer Overflow. In this case, the model successfully identifies the vulnerability type, leveraging the high quality of the decompiled program (ED equal to 88%).
4.4. Impact of decompilation quality on vulnerability prediction
The proposed solution leverages the output of the neural decompilation as the input for the vulnerability prediction. Consequently, it is both valuable and insightful to investigate how the quality of the decompiled code influences the performance of the subsequent classification model.
Table 8. Qualitative analysis on multi-class classification.
Example 1 Example 2
Decompiled C/C++ void funct 0 () { int data ; myUnion ; data = - 1 ; fscanf ( stdin , " ", & data ); myUnion . unionFirst = data ; { int data = myUnion . unionSecond ; { short shortData = ( short ) data ; printShortLine ( shortData ); } } } static int staticFive = 5 ; void funct 0 () { unsigned int data ; data = 0 ; if ( staticFive == 5 ) { data = (unsigned int) RAND 3 2 (); } if ( staticFive == 5 ) { { unsigned int result = data - 1 ; printUnsignedLine ( result ); } } }
ED 70% 88%
CWE observed CWE 369: Divide by Zero CWE 190: Integer Overflow
CWE predicted CWE 197: Numeric Truncation (misprediction) CWE 190: Integer Overflow (correct prediction)
Figure 3. Bar Chart showing the impact of neural decompilation quality (ED-Mean) on vulnerability prediction (F1-Score) for each classified CWE.
Figure 3 illustrates the comparison between the F1-Score and ED-Mean (Edit Distance Mean) across the classified CWE categories. It is worth noting that CWEs as 762, 36, and 23 present high values for both ED-Mean F1-Score. In contrast, CWE 197 presents extremely low values for both F1-Score and ED-Mean. To investigate the relationship between the models’ performance on these tasks we analyze the correlation between the two phenomena. We adopted Pearson’s correlation coefficient, denoted as $r$, with $x$ as the F1-Score, $y$ as the ED-Mean, and $n$ as the number of $(x,y)$ pairs (i.e., the number of CWE entries in this case). The computed correlation coefficient between the F1-Score and the ED-Mean is:
$r=\frac{n\left(\sum xy\right)-\left(\sum x\right)\left(\sum y\right)}{\sqrt{\left[n\sum x^{2}-\left(\sum x\right)^{2}\right]\left[n\sum y^{2}-\left(\sum y\right)^{2}\right]}}=\textbf{0.65}$
The result indicates a moderate, positive correlation. This suggests that as the ED-Mean improves, the average F1-Score value tends to increase correspondingly, reflecting a relationship where higher classification performance is associated with smaller error distances.
On the reverse, as demonstrated in the first example in Table 4.4, it is still plausible that a high ED score may lead to a low F1-Score. Even though the translation may present a high quality standard, two kinds of issues may still arise: i) the decompiled code may be logically flawed, ii) the vulnerability prediction process may have failed to identify the underlying vulnerability.
5. Lessons Learned
The results of the experimental study provided us with several lessons learned about using neural decompilation in the cybersecurity landscape.
Applicability of neural decompilation to cybersecurity. The idea of leveraging neural decompilation for cybersecurity-related tasks is promising. Specifically, given the quality of the translated code, neural decompilation showed great potential for vulnerability prediction, with the possibility to unlock several novel applications, e.g., binary malware recognition. We analyzed three state-of-the-art models, i.e., fairseq, CodeBERT, and CodeT5+, with fairseq emerging as the best-performing model, achieving quality translations comparable to other studies in code translation and providing a representative basis for our experiments. Future work includes integrating additional models for neural decompilation, with a deeper inspection of pre-trained NMT models.
Use of DL for vulnerability prediction. The results of the vulnerability prediction task revealed that DL models were indeed appropriate for vulnerability detection and identification, with transformer-based models emerging as the best performing. The selected models could accurately predict whether a C/C++ function was vulnerable and the related vulnerability in the MITRE CWE taxonomy.
@@ 6. Related Work
Neural decompilation is a recent and innovative research trend in binary analysis. Hosseini et al. (hosseini2022beyond, ) introduced Beyond The C, a neural decompiler based on the fairseq model. This solution is retargetable for several high-level programming languages such as C, OCaml, Go, and Fortran. The results of their study are promising for all languages, scoring the best result in terms of ED ($54\%$) on the decompilation of C/C++ code (training on $2,000,000$ samples). Katz et al. (katz2018using, ) proposed an approach for neural decompilation using RNNs. The solution was trained and evaluated on snippets of binary machine code compiled from C source code. This solution achieved an ED equal to $30\%$, leveraging about $700,000$ samples.
The neural decompilation phase of our experimental study builds on top of the solution proposed by Hosseini et al. but differently, we selected a different set of programs and performed a comparative analysis between state-of-the-art NMT models. In addition, rather than focusing on neural decompilation only, our study is focused on applying this novel solution to the cybersecurity landscape, in particular vulnerability prediction on binary code.
@@
Zheng et al. (zheng2019recurrent, ) devised an approach to identify vulnerabilities in binary files using deep learning (DL) models, in particular Recurrent Neural Networks (RNN). To this end, this work leverages the Juliet C/C++ dataset (Juliet, ), described in Section 2.2. Then, the binary code is lifted to LLVM (LLVM, ) intermediate representation (IR), constituting the input for the RNNs. This solution can identify 4 CWEs with 90% accuracy. Schaad et al. (schaad2023deep, ) extended the previous approach (zheng2019recurrent, ), leveraging several DL models, i.e., Recurrent Neural Networks (RNN), Long-Short Term Memory (LSTM), and Gated Recurrent Unit (GRU). Their solution can identify if a binary file is vulnerable with 88% accuracy, and classify 23 CWEs out of the 118 described in the Juliet dataset with 77% accuracy. Pewny et al. (pewny2015cross, ) designed a cross-architecture solution to identify bug signatures in binary executables: the binaries are lifted to IR to unify the different assembly languages and then converted to control flow graphs (CFG) to compare basic blocks to the known signatures. This solution can identify 5 technology-specific vulnerabilities across three architectures (ARM, MIPS, x86), including the well-known Heartbleed vulnerability for OpenSSL (HeartbleedCVE, ; Heartbleed, ). DiscovRE (eschweiler2016discovre, ) is a prototype for identifying binary code bugs. It can identify two vulnerabilities, i.e., Heartbleed and POODLE for OpenSSL and SSLv3 respectively, and supports four architectures, namely x86, x64, ARM, and MIPS. DiscovRE measures structural similarities between the target code and the vulnerable prototype through custom numeric features and metrics to identify the vulnerabilities. Dahl et al. (dahl2020stack, ) developed a solution to identify stack-based buffer overflow vulnerabilities in binary files, along with a custom dataset for this task. First, the binary code is disassembled and then compressed to discard the redundant information. Again, the vulnerability detection model is an LSTM, which achieves 90% accuracy in detecting stack-based buffer overflows. VDISCOVER (grieco2016toward, ) is a ML-based tool for memory corruption detection. It comes with a custom dataset, VDISCOVERY, built from 1,039 programs of the Debian bug tracker. VDISCOVER can identify 4 types of memory corruptions, namely heap memory, stack memory, corruption of argument in a function, and corruption of a pointer to a function, with 55% accuracy. Table 7 summarizes the comparison between the related work and our solution. Our work differs from the previous ones, in terms of binary code analysis and vulnerability prediction. On the one hand, neural decompilation does not need any formal representation of the binary code (e.g., control flows, intermediate representation) since the code is treated as plain text for the decompilation task. The use of NMT for this task reduces the complexity of the decompilation process, requires less knowledge about the source and target languages, and makes the solution language independent, i.e., it is potentially applicable to a different programming language or an alternative target architecture with the only cost of re-training the model. Indeed, we can observe that only two works, namely DiscovRE (eschweiler2016discovre, ) and the work from Pewny et al. (pewny2015cross, ), are retargetable for the architecture. On the other hand, contrary to other works that focus on technology-specific or single vulnerabilities, we focus on identifying several vulnerabilities described in standard taxonomies, namely the CWE provided by MITRE, achieving higher accuracy than two related works (pewny2015cross, ; grieco2016toward, ) and also surpasses the only work that encompasses a comparable number of vulnerabilities (schaad2023deep, ). Even though the other works achieve a slightly better performance, they only analyze a significantly smaller number of technology-specific vulnerabilities (zheng2019recurrent, ; eschweiler2016discovre, ; dahl2020stack, ). Using a standardized vulnerability taxonomy is a key aspect in evaluating these works: technology-specific vulnerabilities are harder to generalize and typically do not apply to different systems, while standardized vulnerabilities, e.g., CWEs, represent global definitions that span several kinds of software systems and products.
7. Threats to Validity
Dataset. The chosen dataset, Juliet C/C++ 1.3, is a synthetic dataset, i.e., novel data created from real data with the same statistical properties. Using a synthetic dataset provides a large amount of samples, accurately annotated with ground truth. However, the synthetic data do not perfectly align with actual data regarding the size and complexity of real code. Therefore, the vulnerability detection accuracy on synthetic data may be higher than in real data. This may hinder the external validity of the study. Nevertheless, the synthetic data still allow us to analyze the decompilation task and the prediction of vulnerabilities in the translated code, which represents a good starting point for this research direction. Moreover, it is challenging to find datasets in the literature that satisfy the requirements for vulnerability prediction in binary code, as illustrated in Section 2.2. Specifically, Juliet represents a fair trade-off between the requirements for the dataset: i) a dataset composed of compilable C/C++ functions and ii) labeled with the vulnerabilities from a standardized taxonomy.
Models selection. The selection of transformer-based and DL models for the neural decompilation and vulnerability prediction tasks may impact the study’s external validity. On the one hand, we selected three state-of-the-art models, namely fairseq, CodeBERT, and CodeT5+, for neural decompilation performing a comparative study to assess their capabilities for this task. On the other hand, we leveraged CodeBERT, CodeT5+, CodeGPT, and traditional DL models for vulnerability prediction since they represent the most popular solution for this task (zheng2019recurrent, ; schaad2023deep, ; dahl2020stack, ; chen2023diversevul, ). We tried to include, to the best of our knowledge, models with different architectures and capabilities, representative of the state-of-the-art.
Metrics. To evaluate the performance of neural decompilation, we selected output similarity metrics such as BLEU, METEOR, ED, and ROUGE. This choice may threaten the construct validity since such metrics may not capture the correctness of the decompiled code. Nevertheless, output similarity metrics represent the most popular practice for evaluating NMT models, both for code-to-code (as in similar works on neural decompilation (hosseini2022beyond, ; katz2018using, )) and NL-to-code translation tasks (liguori2023evaluates, ). Specifically, ED represents the prevalent metric to estimate the performance of code-to-code translation.

### v2

@@ 4.1. Evaluation - Neural Decompilation
We evaluated the translation quality using output similarity metrics including BLEU, METEOR, ROUGE, and Edit Distance (ED), illustrated in Table 2. Each metric expresses different characteristics of translation quality. One of the most relevant metrics in code translation is ED, suitable for capturing small variations in the code and representative of the number of changes to perform on the decompiled output. We refer to this metric as ED for consistency with prior work, but it represents a normalized ED measure (Edit Similarity) ranging from 0, the worst case, to 1, the best case.
@@
Ground Truth Neural Decompilation Ghidra
static void funct0(long *data) { static char* funct0(char *data){ void funct0(long *param_1){
delete data; data = new char[100]; if (param_1!=(long *)0x0) {
} return data; operator.delete(param_1,8);}
void funct1(){ } return;
long * data; void funct1(){ }
data = NULL; char * data ; void funct1(void) {
data = new long[100]; data = NULL ; long *plVar1;
funct0(data); data = funct0(data); plVar1 = (long *)operator.new[](800);
} delete data ; funct0(plVar1);
} return;
ED Score 57% 37%
Table 2 compares different decompilation solutions based on their approach, number of functions used for training and testing, maximum target length, and output similarity metrics. The results of neural decompilation are outstanding. Specifically, fairseq achieves a 59% ED score using 75,1497514975,14975 , 149 C/C++ functions for both training and testing with a maximum target length of 338 tokens. These results are close to those of other studies on this topic: Hosseini et al. (hosseini2022beyond, ) reached an 54% ED using about 2,000,00020000002,000,0002 , 000 , 000 C/C++ functions with a maximum length of $271271$, while Katz et al. (katz2018using, ) achieves an ED of 30% using about 700,000700000700,000700 , 000 C-code snippets with a maximum target length of $8888$. It is worth noting that both studies leverage a number of samples at least an order of magnitude higher than ours. In addition, the function lengths analyzed with our approach are greater than those used by either work. For the sake of clarity, we note that these works adopt a complementary version of ED, less widespread for NMT tasks (liguori2023evaluates, ). Focusing on the other metrics obtained from fairseq, ROUGE-L shows that, in 77% of cases, the generated code contains word sequences similar to those of the reference. We observe a 4-gram overlap of 58% (BLEU-4) and a token order adherence of 64% (METEOR).
For the sake of completeness, we also evaluated the performance of a traditional decompiler, Ghidra (Ghidra, ). Specifically, we randomly sampled 1,000 programs from our decompiled test set. Ghidra relies on control flow graphs (CFGs), data flow graphs (DFGs), and other intermediate representations (IRs) to perform decompilation. Ghidra results, with a 23% ED, indicate significantly lower performance compared to studies utilizing neural network-based models, which suggests that traditional graph-based methods may struggle to capture the complexities of the source code as effectively as neural approaches. However, unlike a traditional decompiler, an NMT model always generates an answer despite having a low probability of approaching the correct one. In addition, one of the disadvantages of using a neural network is that there is less control over the translation, i.e., in case of errors it is very complex to determine which factor led the model to a specific decision over others.
Table 4. Models performance for the binary and multi-class classification task. Best performance is blue.
Binary Classification Multi-class Classification
Approach Model Accuracy (%) F1-Score (%) Accuracy (%) F1-Score (%)
Flawfinder* (flawfinder, ) Static Analysis 52 49 13 11
Schaad et al.° (schaad2023deep, ) SRNN 88 88 77 79
Proposed approach SRNN 88 88 72 70
LSTM 94 94 75 72
GRU 91 90 78 77
CodeT5+ 94 94 81 81
CodeGPT 95 95 82 82
CodeBERT 94 94 83 82
@@
*: multi-class concern Flawfinder-compatible CWEs.        °: multi-class concern different CWEs selected by Schaad et al. (schaad2023deep, ).
An interesting finding is that pre-trained models, i.e., CodeT5+ and CodeBERT, perform considerably worse than the model trained from scratch, i.e., fairseq. As noted by Tufano et al. (tufano2023automating, ), we can associate the negative impact of pre-training with the size of the training dataset. Pre-training helps boost performance only if the amount of fine-tuning data available is small. We deem our dataset with about 75,000 samples large enough to overwrite the parameters tuned in the pre-training phase. According to these results, we adopt fairseq as the neural decompiler for further analysis.
Qualitative analysis. Table 3 illustrates an example of ASM-C/C++ code translation. In the reported vulnerable example associated with CWE 762 Mismatched Memory Management Routines, the potential flaw is to deallocate the array data in heap memory with delete instead of delete []. Specifically, we report two versions of decompiled code for the example, one produced by our neural decompilation model (fairseq) and the other through Ghidra (Ghidra, ).
On the one hand, it is possible to notice that neural decompilation retraces the original code structure. The resulting code achieves an ED score of 57%. Although some operations are reversed in the corpus of the two functions, the logical sense and semantics of the code are correct. In addition, the code is extremely readable. However, this is not always the case. Sometimes, the translated program leads us to consider that although ED may be high, the logical sense of the program could be incorrect.
On the other hand, the code produced by Ghidra closely resembles the assembly structure and scores a 37% ED. It is worth noting that this code is harder to understand for a human operator. Moreover, to perform decompilation, Ghidra needs the entire assembly file, not just the function to decompile. However, the key advantage of Ghidra, as a traditional decompiler, is that it offers a higher probability that the decompiled code is functionally and semantically correct, though this is not guaranteed.
@@
Work Dataset Classes of Standardized Intermediate Vulnerability Bi-Classification Multi-Classification
Vulnerabilities Taxonomy Representation Predictor Accuracy Accuracy
@@
4.2. Evaluation - Vulnerability Prediction
Binary Classification. The first analysis concerns the classification of the target code as vulnerable/non-vulnerable. We submitted the C/C++ code translated by the neural decompilation model to classification, adopting the same data pre-processing for each model. We split the data 80-20%, ensuring stratification by target and maintaining a 50-50% balance between vulnerable and non-vulnerable samples. We selected accuracy and F1-Score as the metrics to evaluate the performance of the models.
Table 4 shows the models’ performance for the binary classification on the decompiled C/C++ code. CodeGPT is the best-performing model with an accuracy and F1-Score of 95%. The highest gap lies between SRNN and CodeGPT, exhibiting a discrepancy of about 7% in terms of F1-Score. Table 4 also reports the performance of a state-of-the-art work on vulnerability prediction in binary files (schaad2023deep, ), which achieves an 88% F1-Score through the LLVM IR and the SRNN model, further described in Table 5 and Section 5.
Furthermore, we tested a static analysis tool, Flawfinder (flawfinder, ), on the same test set. This tool scans C/C++ source code and flags potential security holes. Flawfinder exhibits metrics that are close to 50% on average, which is equivalent to a random model. Therefore, the results support the implemented design choice of using DL models for vulnerability prediction instead of a static analysis tool. Finally, we leveraged Shap (NIPS2017_7062, ), a tool for explaining the output of our vulnerability prediction models, to better understand which features are the most influential for the learning process, i.e., the tokens most exploited for classification. This is a helpful analysis to avoid falling into the so-called Clever Hans effect, which occurs when the model produces correct predictions based on ”wrong” features (kauffmann2020clever, ). The results revealed tokens in C/C++ code such as: if, new, strcat, funct, static, malloc, and source. These results allow us to state that the model focuses on tokens that are relevant to implementing countermeasures to vulnerabilities, e.g., if constructs to check specific conditions (e.g., underflow or overflow).
Multi-class Classification. This analysis aims to identify the CWE associated with a vulnerable function. Specifically, the number of classes (CWEs) for this classification task is 20 (190, 762, 121, 590, 194, 197, 78, 36, 23, 690, 369, 457, 134, 680, 758, 400, 789, 253, 398, 761), plus an additional class associated with non-vulnerable (good) functions. We obtain a total of 42,483 samples, with 20,393 corresponding to vulnerable functions and 22,090 to non-vulnerable ones. The complexity of this task stems from similarities between test cases across classes. Rather than selecting classes solely by frequency, we manually chose those with distinct CWEs to minimize model confusion with similar ones. Consequently, to the best of our knowledge, we evaluated the difference between CWEs from a logical point of view. We prioritized the CWEs that belonged to the Top 25 CWEs published yearly by MITRE among the most represented ones, i.e., those with a reasonable amount of samples. Finally, we split the data 80%-20%, ensuring stratification by CWE.
@@ 4.2. Evaluation - Vulnerability Prediction
Table 4 shows the accuracy and F1-Score of the models for the multi-class classification task. The models’ performance decreased from the binary classification case, with CodeBERT performing the best with an accuracy of 83% and F1-Score of 82%. The performance gap between our approach, utilizing CodeBERT, and the method proposed by Schaad et al. (schaad2023deep, ) is approximately 6% in terms of accuracy and 3% in the F1-score. However, it is important to note that the 20 classes analyzed in our study differ both in type and quantity from their 23 classes (schaad2023deep, ), as shown in Table 5.
@@ 5. Related Work
Table 5 summarizes the comparison between the related work and our solution. Our work differs from the others, in terms of binary code analysis and vulnerability prediction. On the one hand, neural decompilation does not need any formal representation of the binary code (e.g., control flows, intermediate representation) since the code is treated as plain text for the decompilation task. On the other hand, contrary to other works that focus on technology-specific or single vulnerabilities, we focus on identifying several vulnerabilities described in standard taxonomies, namely the CWE provided by MITRE, achieving higher accuracy than two related works (pewny2015cross, ; grieco2016toward, ) and also surpasses the only work that encompasses a comparable number of vulnerabilities (schaad2023deep, ). Even though the other works achieve a slightly better performance, they only analyze a significantly smaller number of technology-specific vulnerabilities (zheng2019recurrent, ; eschweiler2016discovre, ; dahl2020stack, ). Using a standardized vulnerability taxonomy is a key aspect in evaluating these works: technology-specific vulnerabilities are harder to generalize and typically do not apply to different systems, while standardized vulnerabilities, e.g., CWEs, represent global definitions that span several kinds of software systems and products.